Mars and Earth orbit the Sun at radii of 228 million km and 150 million km, respectively. When, in the future, your friend from Mars calls you on the phone and you answer "Hello," what are the minimum and maximum times you will have to wait for your friend to reply?
Answer #1

Minimum separation = (228 - 150) million km = 78 million km = 7.8E10 m
Maximum separation = (228 + 150) million km = 378 million km = 3.78E11 m

Minimum time = (7.8E10 m)/(3.00E8 m/s) = 2.6E2 s = 260 s = 4.3 minutes
Maximum time = (3.78E11 m)/(3.00E8 m/s) = 1.26E3 s = 1260 s = 21.0 minutes